Shift Summary Evening Shift Nov 5

B. Cox (cox@fnal.gov)
Sat, 6 Nov 1999 00:15:38 -0600 (CST)

Shift Summary for Evening Shift, Nov. 5/99

On Shift: Brad C., Sydney T.

Hours of beam available at all intensities: 6.5 hours
Hours of beam available at nominal intensity: 5.5 hours
Hours of E799 data (special runs not included): 4.7 hours
Hours of special runs (laser scans, muon runs...): 0 hours

Accounting of lost time:

tuning of beam to 9E12 produces sporatic beam ~45 minutes
DDD veto 3/KTeV1 reboot ~hour and 10 minutes
TeV quenches ~hour and 34 minutes
Beam aborts ~10 minutes
Pipeline problems ~10 minutes

========================================================================
Runs ended this shift:

14911 BEAM_799
14913 BEAM_799

=======================================================================
Tapes taken to FCC this shift

none

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

15:30 Begin shift. Waiting for beam

16:10 Beam begins to come back. Tuning toward 9E12

16:25 DDD veto 3 error. DDD3 stops working. We start a new run to try to
get it back

16:30 Try a fresh start, KTeV1 is not responding. Cannot log into it.
Need to reboot apparently. Contact experts to get the root
directory password. Seems not to work. Spend some time trying
to break the code.

16:35 Beam off

16:45 Beam back......on and off

16:55 beam drifts up to over 1E13. Call MC to ask them to lower to
9E12

17:00 RF drops by and proclaims the beam tuned and autotune functioning
properly,

17:07 Finally figure out the password after trying 10**6 combinations..
proceed to reboot.

17:25 Reboot of KTeV1 finished. Try to start run.

17:43 Run 14913 started;

18:22 CsI 1163 gain was wrong. Run stopped and gain fixed.

18:26 TeV quench

18:31 Start run 14914 BEAM_799

19:00 beam returns

19:05 TeV quench

20:00 Beam returns; MC thinks they know what causes the TeV quenches but
they will start slow a low intensity for a while ~ 3E12

20:55 KTEVANA errors - missing FFFF word, stream 5, RWP:9.

22:45 List of "minor" problems listed in Log xxii on page 167.
appropriate experts shoudl take a look.

22:50 abort,no beam

23:00 beam returns

23:20 Pipeline errors; resetting pipeline.
Hard reboot of pipeline crate 3
Lost permit in stream 5 recovered by power-cycling repeater in
laser crate.

23:30 Back running

========================================================================
Description of Shift:

========================================================================
Follow up Notes on Previous Shift